# GPU-Accelerated Routability-Driven Automated Placement for Large-Scale Photonic Integrated Circuits

### v1

@@ Abstract.
As technology advances, photonic integrated circuits (PICs) are rapidly scaling in size and complexity, with modern designs integrating thousands of components to meet the demands of artificial intelligence (AI), high-performance computing, and chip-to-chip optical interconnects. However, the analog custom layout nature of photonics, the curvy waveguide structures, and single-layer routing resources impose stringent physical constraints, such as minimum bend radii and waveguide crossing penalties, which make manual layout the de facto standard. This manual process takes weeks to complete and is error-prone, which is fundamentally unscalable for large-scale PIC systems. Existing automation solutions have adopted force-directed placement on small benchmarks with tens of components, with limited routability and scalability. To fill this fundamental gap in the electronic-photonic design automation (EPDA) toolchain, we present Apollo, the first GPU-accelerated, routing-informed placement framework tailored for large-scale PICs. Apollo features an asymmetric bending-aware wirelength function with explicit modeling of waveguide routing congestion and crossings to preserve enough routing spacing for routability maximization. Meanwhile, conditional projection is employed to gradually enforce a variety of user-defined layout constraints, including alignment, spacing, etc. This constrained optimization is accelerated and stabilized by a custom blockwise adaptive Nesterov-accelerated optimizer, ensuring stable and high-quality convergence. To catalyze research in PIC layout automation, we also develop and open-source large-scale PIC benchmarks derived from real-world photonic tensor core designs. Compared to existing methods, Apollo can generate high-quality layouts for large-scale PICs with an average routing success rate of 94.79% across all benchmarks within minutes. By tightly coupling placement with physical-aware routing, Apollo establishes a new paradigm for automated PIC design—bringing intelligent, scalable layout synthesis to the forefront of next-generation EPDA. We will open-source our code.
@@ 1. Introduction
Despite their immense potential, the physical layout of PICs remains largely a manual and iterative process, even in state-of-the-art commercial and academic design flows (14). Unlike digital VLSI, where placement and routing are decoupled and highly automated, PIC layout presents fundamentally different challenges due to the analog nature of light propagation and the unique physical constraints needed to maintain signal integrity, performance, and fabrication yield. In PICs, we claim that waveguide routing must be tightly integrated into the placement process, far more so than in digital VLSI, because the feasibility and quality of routing are highly sensitive to component placement. At the core of these challenges lies the fact that photonic waveguides are directional, curvilinear, and highly space-consuming (29). Unlike metal wires in VLSI circuits or on PCBs, waveguides must conform to minimum bend radii to avoid excessive propagation loss and cannot intersect with photonic devices. Waveguide crossings, analogous to vias in VLSI, must also be carefully planned to minimize crosstalk, insertion loss, and excessive detours, all of which are tightly influenced by placement. Crossings further introduce non-negligible area overhead and impose strict port orientation requirements, complicating the layout even more. Moreover, PICs are typically restricted to a single layer (or a few for advanced PICs), resulting in severe layout resource contention. These properties make waveguide routing a dominant constraint during placement. Furthermore, PICs often involve mixed-size components, ranging from compact 2$\times$2 $\mu m^{2}$ devices (e.g., micro-disk modulators and y-branches) to large 100$\times$1000 $\mu m^{2}$ devices (e.g., Mach-Zehnder modulators, MZM). Some devices exhibit extreme aspect ratios (e.g., 1:20), posing additional challenges for placement that must handle geometric diversity, routability, and photonic-specific layout rules.
@@ 2.1. Photonic Circuit Placement
Commercial PIC physical design toolflow has traditionally relied on schematic-driven layout methodology (14), in which devices are placed according to their logical positions and then manually abutted via waveguide ports, a time-consuming process that does not scale to large systems. To boost productivity, the work (9) introduced a semi-automatic Photonic CAD tool for visual place-and-route of on-chip photonic networks. Fully automatic frameworks such as Proton (2) and Platon (24) then focused on minimizing insertion loss, particularly at crossings, while another work (6) further reduced loss through device flipping and rotation. A recent research incorporated critical-path insertion-loss optimization directly into placement (5). Another work (23) jointly optimized network topology and its physical realization. Thermal-aware approaches have also been proposed (11). However, these efforts largely overlook the geometric cost of waveguide routing itself. In this work, we argue that a routing-informed placement engine, one that explicitly accounts for bends, crossings, and congestion during device placement, is essential for truly scalable, routable PIC layout.
@@ 3.1. Understanding the Placement Challenges of Photonic Circuits
PICs are limited to one or very few routing layers. This not only leads to scarce routing resources but also makes 90-degree waveguide crossings topologically inevitable in many designs, introducing insertion loss, crosstalk, and significant area overhead (as crossings take spaces like via in VLSI/PCB), further increasing congestion and routability complexity.
@@
Mixed-size PIC components are geometrically diverse. Photonic components span from compact 2$\times$2 $\mu m^{2}$ compact splitters/resonators to millimeter-scale high-speed modulators with extreme aspect ratios (e.g., 1:20). This diversity mirrors the challenges of mixed-size VLSI placement, long known to be combinatorial and ill-conditioned for joint optimization.
Lack of high-quality PIC routers. In PIC layout, placement quality is tightly coupled with the capabilities of the router. However, the current EPDA ecosystem lacks mature, high-quality routers for PICs. This practical limitation means that many routing complexities (15), such as bend management, crossing avoidance, and port alignment, must be preemptively handled during placement. As a result, the burden of achieving a routable and high-performance layout is currently disproportionately shifted to the placement engine.
@@ 3.3. PIC-Specific Placement Objective Functions
Weighted-average wirelength (WA-WL) is widely used in VLSI placement for wirelength cost, which is used to approximate the half-perimeter wirelength (HPWL). However, unlike VLSI routing, PICs only have two-pin nets to describe waveguides and use ports instead of pins to connect the waveguides. These ports are essentially open-ended waveguide segments that can only be accessed in a specific orientation. Therefore, HPWL is not an appropriate approximation for the PIC waveguide routes. The original WA-WL is symmetric along the x and y-axis. However, as shown in Fig. 1, PIC does not prefer sharp turns. Instead, we need to use a smooth curvy waveguide for bending, which is area-consuming and thus harms routability.
@@
$\cos\theta_{i}$ $=\frac{\mathbf{w}\cdot\mathbf{v}}{\|\mathbf{w}\|\|\mathbf{v}\|},\quad i=1,2$
@@
$\bm{w}$ is defined as the vector from the first pin to the second pin of net and $\bm{v}$ is the unit vector indicating the port orientation. Then $\theta_{i}$ represents the angle between the port orientation and straight line connection of two-pin nets, as shown in Fig. 2. In this way, a large $\theta$ (e.g., 150-degree) will result in a large penalty factor $W_{\theta}$, so that we can penalize the net with a large bending angle. $(\cdot)_{+}$ is the ReLU function that casts a negative value to 0, which makes sure no extra penalty is added for acute angles. The parameter $c\in[0,1]$ is used to control the angle margin. The $\mathrm{cosWA}$ will become 0 when both $\theta_{1}$ and $\theta_{2}$ are satisfied. As shown in Fig. 2, the $\mathrm{cosWA}$ function is asymmetrical along the x-axis if the corresponding port has a horizontal orientation. Note that y-axis cosWA wirelength in this case will degrade to a symmetric WA wirelength. Similarly, it simply flips to x-symmetry and y-asymmetry for vertical ports.
Another key variable here is the exponent $\alpha$ that makes this function not an approximation of linear HPWL, which leads to very long wires and is unable to encourage symmetric layouts as many placement solutions can have the same HPWL. To borrow the long-wire penalty property of quadratic wirelength without overestimating long waveguides and ignoring short waveguides, we add this $\alpha\in[1,2]$ exponent to make it smoother. Empirically, we set $\alpha$ to 1.4, which is later identified as the optimal setting in the ablation study section.
Figure 3. Proposed bending-aware wirelength function: detour region results in larger wirelength cost
@@
A major root cause of routing failure in PICs is spacing deficiency. Unlike VLSI, where multiple metal layers are available, PICs are typically limited to a single optical routing layer. As a result, routing resources are scarce, and waveguide crossings and curved bends consume significant area. Without enough spacing, routing is likely to fail (see Fig. 3(a)). Hence, it is essential to integrate routing-informed spacing estimation into the placement process. Accurate modeling of the spatial demands induced by port density, bend radii, and anticipated crossings can preempt routing congestion and ensure sufficient space for successful waveguide routing.
Informed by empirical analysis of the routing behavior for the open-source automated PIC router LiDAR, the main routing failure comes from the port accessibility and waveguide crossing insertion, we propose the routing-informed net spacing model $S_{i}$ based on the port density $P_{\text{dens}}$ and routing congestion $R_{\text{cong}}$. For each net, we define a spacing demand based on two factors: (1) the local port density, which captures how many co-directional ports are clustered together, and (2) estimated routing congestion, which accounts for potential waveguide crossings. Specially, for a 2-pin net $i$ connecting port $p_{1}$ (on node $v_{1}$) and port $p_{2}$ (on node $v_{2}$), the net spacing demand $S_{i}$ is estimated as
@@
$P_{\text{num}}(v,p)$ =|$=\big{|}\{p_{j}\in\text{Ports of }v~{}~{}|~{}~{}\text{Dir}(p_{j})=\text{Dir}(p)\}\big{|}.$
where the local port density $P_{\text{dens}}$ represents the basic spacing needed for feasible port access, as shown in Fig. 3(b). The $r_{\text{bend}}$ is the bending radius, e.g., 5 $\mu m$. $P_{\text{num}}(v,p)$ is the number of ports on node $v$ that have the same orientation as port $p$ connected to net $i$. $S_{\text{crs}}$ is the waveguide crossing height/width (typically a 4-port square-shape device), e.g., 10 $\mu m$. Since a net connects two ports of different components, which gives different port density spacing for the net, we choose the one with the largest required spacing.
Apart from the basic spacing $P_{\text{dens}}$, net crossings will also cost a large amount of routing resources that further cause congestion issues, which should be considered in the spacing model. Once the positions of all components have roughly stabilized, we estimate the potential crossings to calculate the additional spacing required. We estimate the crossings by calculating the line segment intersections between nets during placement. If a crossing occurs, we approximate the additional spacing needed for cascaded crossings (see Fig. 3(c)) by multiplying the number of estimated crossings #$\#CR_{net}$ by the size of a crossing $S_{\text{crs}}$,
@@
Since net crossings #$\#CR_{net}$ evolve as components move, the congestion estimate is periodically updated to improve accuracy, e.g., every 100 iterations.
Finally, we define a spacing penalty term in the objective function,
(4) $NS(x,y)=[(v_{x}\cdot dx-S_{x})_{+}]^{2}+[(v_{y}\cdot dy-S_{y})_{+}]^{2},$
@@
The following formula is the combined placement objective function used in Apollo:
(6) $\mathcal{L}(x,y)=\textrm{cosWA}(x,y)+\lambda_{NS}\cdot NS(x,y)+\mathcal{D}(x,y),$
where $\lambda_{NS}$ is the weight of the spacing penalty which is set as 1 in the later experiment. At the beginning of the placement process, all components are randomly positioned at the center of the layout, with all the components overlapping each other. Then, driven by wirelength and density, each component will move to its appropriate position, while the spacing penalty will force components in congested areas to separate from each other, satisfying the spacing requirements.
@@ 3.4. Designer-Constrained Placement via Conditional Projection and Cell Inflation
(10) $x_{\text{new},i}=x_{\min}+\text{round}\left(\frac{(|G|-1)(x_{i}-x_{\min})}{x_{\max}-x_{\min}}\right).$
As optimization progresses, the increasing sharpness $s_{t}$ ensures that constraint satisfaction is gradually enforced, allowing initial placement flexibility while converging smoothly to the final aligned or evenly spaced configuration. Note that this projection is not performed after the optimizer descent step. Otherwise, it is not aware of this projection. The projection is applied after the tentative descent used for step size estimation in the Nesterov optimizer (see Alg. 1 Line 18).
@@ 3.5. Stabilizing Mixed-Size PIC Placement via Blockwise Adaptive Nesterov Optimizer
As emphasized in Section 3.1, the significant heterogeneity in component sizes presents a major challenge for stability and convergence in photonic IC placement. To address this, we introduce a Blockwise-Adaptive Barzilai–Borwein (BBB) step size scheme in Nesterov-accelerated gradient descent optimizer in Alg. 1, named BNAG, specifically tailored for mixed-size PIC layouts. In our formulation, all placement variables are partitioned into 4 logical blocks: movable instances and dummy fillers in both the $x$ and $y$ dimensions. An independent Barzilai–Borwein (3) step size is computed for each block, and the resulting values are truncated to prevent instability. This blockwise treatment allows us to stabilize the motion of large nodes by decoupling their updates from smaller cells and encourages faster convergence of filler cells to surround movable instances and fill their spacing. While the Barzilai–Borwein (or inverse Lipschitz) step size is effective for acceleration, it often overestimates the appropriate step size, especially when certain cells are already near-optimal. To mitigate this and ensure convergence, we apply a global cosine annealing schedule to gradually reduce the effective step size over time. This helps maintain robustness in the presence of highly uneven cell dimensions and density.
@@
1:$a_{k},\;u_{k}(\text{curr. sol.}),\;v_{k},\;v_{k-1},\;\nabla f(v_{k}),\;\nabla f(v_{k-1}),\;\{B_{1},\dots,B_{m}\}$
2:$u_{k+1}(\text{updated sol.}),\;v_{k+1},\;a_{k+1}$
@@
17:$a_{k+1}\leftarrow\frac{1+\sqrt{4a_{k}^{2}+1}}{2}$
18:$v_{k+1}\leftarrow u_{k+1}+\frac{a_{k}-1}{a_{k+1}}(u_{k+1}-u_{k})$
19:$\eta_{k+1}\leftarrow\eta_{\min}+\frac{1}{2}(\eta_{0}-\eta_{\min})\left(1+\cos\left(\frac{\pi k}{K_{\max}}\right)\right)$
20:project$(v_{k+1})\quad\triangleright$ enforce constraints

### v2

# Apollo: Automated Routing-Informed Placement for Large-Scale Photonic Integrated Circuits

@@ Abstract.
As technology advances, photonic integrated circuits (PICs) are rapidly scaling in size and complexity, with modern designs integrating thousands of components to meet the demands of artificial intelligence (AI), high-performance computing, and chip-to-chip optical interconnects. However, the analog custom layout nature of photonics, the curvy waveguide structures, and single-layer routing resources impose stringent physical constraints, such as minimum bend radii and waveguide crossing penalties, which make manual layout the de facto standard. This manual process takes weeks to complete and is error-prone, which is fundamentally unscalable for large-scale PIC systems. Existing automation solutions have adopted force-directed placement on small benchmarks with tens of components, with limited routability and scalability. To fill this fundamental gap in the electronic-photonic design automation (EPDA) toolchain, we present Apollo, the first GPU-accelerated, routing-informed placement framework tailored for large-scale PICs. Apollo features an asymmetric bending-aware wirelength function with explicit modeling of waveguide routing congestion and crossings to preserve enough routing spacing for routability maximization. Meanwhile, conditional projection is employed to gradually enforce a variety of user-defined layout constraints, including alignment, spacing, etc. This constrained optimization is accelerated and stabilized by a custom blockwise adaptive Nesterov-accelerated optimizer, ensuring stable and high-quality convergence. To catalyze research in PIC layout automation, we also develop and open-source large-scale PIC benchmarks derived from real-world photonic tensor core designs. Compared to existing methods, Apollo can generate high-quality layouts for large-scale PICs with an average routing success rate of 94.79% across all benchmarks within minutes. By tightly coupling placement with physical-aware routing, Apollo establishes a new paradigm for automated PIC design—bringing intelligent, scalable layout synthesis to the forefront of next-generation EPDA. Our code is open-sourced at link***https://github.com/ScopeX-ASU/Apollo.
@@ 1. Introduction
Despite their immense potential, the physical layout of PICs remains largely a manual and iterative process, even in state-of-the-art commercial and academic design flows (14). Unlike digital VLSI, where placement and routing are decoupled and highly automated, PIC layout presents fundamentally different challenges due to the analog nature of light propagation and the unique physical constraints needed to maintain signal integrity, performance, and fabrication yield. In PICs, we claim that waveguide routing must be tightly integrated into the placement process, far more so than in digital VLSI, because the feasibility and quality of routing are highly sensitive to component placement. At the core of these challenges lies the fact that photonic waveguides are directional, curvilinear, and highly space-consuming (29). Unlike metal wires in VLSI circuits or on PCBs, waveguides must conform to minimum bend radii to avoid excessive propagation loss and cannot intersect with photonic devices. Waveguide crossings, analogous to vias in VLSI, must also be carefully planned to minimize crosstalk, insertion loss, and excessive detours, all of which are tightly influenced by placement. Crossings further introduce non-negligible area overhead and impose strict port orientation requirements, complicating the layout even more. Moreover, PICs are typically restricted to a single layer (or a few for advanced PICs), resulting in severe layout resource contention. These properties make waveguide routing a dominant constraint during placement. Furthermore, PICs often involve mixed-size components, ranging from compact $\times2$ $\mu m^{2}$ devices (e.g., micro-disk modulators and y-branches) to large $\times100$0 $\mu m^{2}$ devices (e.g., Mach-Zehnder modulators, MZM). Some devices exhibit extreme aspect ratios (e.g., 1:20), posing additional challenges for placement that must handle geometric diversity, routability, and photonic-specific layout rules.
@@ 2.1. Photonic Circuit Placement
Commercial PIC physical design toolflow has traditionally relied on schematic‑driven layout methodology (14), in which devices are placed according to their logical positions and then manually abutted via waveguide ports, a time‑consuming process that does not scale to large systems. To boost productivity, the work (9) introduced a semi‑automatic Photonic CAD tool for visual place‑and‑route of on‑chip photonic networks. Fully automatic frameworks such as Proton (2) and Platon (24) then focused on minimizing insertion loss, particularly at crossings, while another work (6) further reduced loss through device flipping and rotation. A recent research incorporated critical‑path insertion‑loss optimization directly into placement (5). Another work (23) jointly optimized network topology and its physical realization. Thermal‑aware approaches have also been proposed (11). However, these efforts largely overlook the geometric cost of waveguide routing itself. In this work, we argue that a routing‑informed placement engine, one that explicitly accounts for bends, crossings, and congestion during device placement, is essential for truly scalable, routable PIC layout.
@@ 3.1. Understanding the Placement Challenges of Photonic Circuits
PICs are limited to very few routing layers. This leads to scarce routing resources and makes 90-degree waveguide crossings topologically inevitable in many designs, introducing insertion loss, crosstalk, and significant area overhead (as crossings take spaces like via in VLSI/PCB), further increasing congestion and routability complexity.
@@
Mixed-size PIC components are geometrically diverse. Photonic components span from compact $\times2$ $\mu m^{2}$ compact splitters/resonators to millimeter-scale high-speed modulators with extreme aspect ratios (e.g., 1:20). This diversity mirrors the challenges of mixed-size VLSI placement, long known to be combinatorial and ill-conditioned for joint optimization.
Lack of high-quality PIC routers. In PIC layout, placement quality is tightly coupled with the capabilities of the router. However, the current EPDA ecosystem lacks mature, high-quality routers for PICs. This practical limitation means that many routing complexities (29; 15), such as bend management, crossing avoidance, and port alignment, must be preemptively handled during placement. As a result, the burden of achieving a routable and high-performance layout is currently disproportionately shifted to the placement engine.
@@ 3.3. PIC-Specific Placement Objective Functions
Weighted-average wirelength (WA-WL) is widely used in VLSI placement for wirelength cost, which is used to approximate the half-perimeter wirelength (HPWL). However, unlike VLSI routing, PICs only have two-pin nets to describe waveguides and use ports instead of pins to connect the waveguides. These ports are essentially open-ended waveguide segments that can only be accessed in a specific orientation. Therefore, HPWL is not an appropriate approximation for the PIC waveguide routes. The original WA-WL is symmetric along the x and y-axis. However, as shown in Fig. 1, PIC does not prefer sharp turns. Instead, we need to use a smooth, curvy waveguide for bending, which is area-consuming and thus harms routability.
@@
$\cos\theta_{i}$ $=\frac{\mathbf{w}\cdot\mathbf{v_{i}}}{\|\mathbf{w}\|\|\mathbf{v_{i}}\|},\quad i=1,2$
@@
Eq. (2) shows the horizontal wirelength of net $e$. $\bm{w}$ is defined as the vector from the first pin to the second pin of the net, where $dx$ and $dy$ represent the horizontal and vertical distances between the two pins, respectively. $\bm{v}$ is the unit vector indicating the port orientation. Then $\theta_{i}$ represents the angle between the port orientation and the straight line connection of two-pin nets. In this way, a large $\theta$ (e.g., 150-degree) will result in a large penalty factor $W_{\theta}$, so that we can penalize the net with a large bending angle as shown in Fig. 2. $(\cdot)_{+}$ is the ReLU function that casts a negative value to 0, which makes sure no extra penalty is added for acute angles. The parameter $c\in[0,1]$ is used to control the angle margin. The $W_{\theta}$ will become 0 when both $\theta_{1}$ and $\theta_{2}$ are within the angle margin defined by $c$. The $\mathrm{cosWA}$ function is asymmetrical along the x-axis if the corresponding port has a horizontal orientation. Note that the y-axis cosWA wirelength in this case will degrade to a symmetric WA wirelength. Similarly, it simply flips to x-symmetry and y-asymmetry for vertical ports.
Another key variable here is the exponent $\alpha$ that makes this function not an approximation of linear HPWL, which leads to very long wires and is unable to encourage symmetric layouts, as many placement solutions can have the same HPWL. To borrow the long-wire penalty property of quadratic wirelength without overestimating long waveguides and ignoring short waveguides, we add this $\alpha\in[1,2]$ exponent to make it smoother. Empirically, we set $\alpha$ to 1.4, which is later identified as the optimal setting in the ablation study section.
Figure 3. Illustration of the critical spacing requirements in PIC routing that should be considered during placement.
@@
A major root cause of routing failure in PICs is spacing deficiency. Unlike VLSI, where multiple metal layers are available, PICs are typically limited to a single optical routing layer. As a result, routing resources are scarce, and waveguide crossings and curved bends consume significant area. Without enough spacing, routing is likely to fail (see Fig. 3(a)). Hence, it is essential to integrate routing-informed spacing estimation into placement. Accurate modeling of the spatial demands from port density, bend radii, and anticipated crossings can preempt routing congestion and ensure sufficient space for successful waveguide routing.
Informed by empirical analysis of the routing behavior for the open-source PIC router LiDAR, the main routing failure comes from the port accessibility and crossing insertion, we propose the routing-informed net spacing model $S_{i}$ based on the port density $P_{\text{dens}}$ and routing congestion $R_{\text{cong}}$. For each net, we define a spacing demand based on two factors: (1) the local port density, which captures how many co-directional ports are clustered, and (2) estimated routing congestion, which accounts for potential waveguide crossings. Specially, for a 2-pin net $i$ connecting port $p_{1}$ (on node $v_{1}$) and port $p_{2}$ (on node $v_{2}$), the net spacing demand $S_{i}$ is estimated as
@@
$P_{\text{num}}(v,p)$ $=\big|\{p_{j}\in\text{Ports of }v~~|~~\text{Dir}(p_{j})=\text{Dir}(p)\}\big|.$
where the local port density $P_{\text{dens}}$ represents the basic spacing needed for feasible port access, as shown in Fig. 3(b). The $r_{\text{bend}}$ is the bending radius, e.g., 5 $\mu m$. $P_{\text{num}}(v,p)$ is the number of ports on node $v$ that have the same orientation as port $p$ connected to net $i$. $S_{\text{crs}}$ is the waveguide crossing height/width (typically a 4-port square-shaped device), e.g., 10 $\mu m$. Since a net connects two ports of different components, which gives different port density spacing for the net, we choose the one with the largest required spacing.
Apart from the basic spacing $P_{\text{dens}}$, net crossings also cost a large amount of routing resources that further cause congestion, which should be considered in the spacing model. Once the positions of all components are roughly stabilized, we estimate the potential crossings to calculate the additional spacing required. We estimate the crossings by calculating the line segment intersections between nets during placement. If a crossing occurs, we approximate the additional spacing needed for cascaded crossings (see Fig. 3(c)) by multiplying the number of estimated crossings $CR_{net}$ by the size of a crossing $S_{\text{crs}}$,
@@
Since net crossings $CR_{net}$ evolve as components move, the congestion estimate is updated periodically every 100 iterations starting from the 100th iteration to improve accuracy. Finally, we define the net spacing penalty for a net $e$ as,
(4) $NS_{e}(x,y)=[(v_{x}\cdot dx-S_{i})_{+}]^{2}+[(v_{y}\cdot dy-S_{i})_{+}]^{2},$
@@
In Apollo, the overall combined objective function is as follows,
(6) $\mathcal{L}(x,y)=\sum_{e\in E}\textrm{cosWA}_{e}(x,y)+\lambda_{NS}\sum_{e\in E}NS_{e}(x,y)+\mathcal{D}(x,y),$
where $E$ denotes the set of all nets, and $\lambda_{NS}$ is the weight of the spacing penalty, which is set to 1 in the later experiment. Initially, all components are randomly placed at the layout center, overlapping each other. Guided by wirelength minimization and density control, components progressively move toward optimized positions, while the spacing penalty pushes apart those in congested areas to meet spacing constraints.
@@ 3.4. Designer-Constrained Placement via Conditional Projection and Cell Inflation
(10) $x_{\text{new},i}=x_{\min}+s_{x}\cdot\text{round}\left(\frac{x_{i}-x_{\min}}{s_{x}}\right),\quad s_{x}=\frac{x_{\max}-x_{\min}}{|G|-1}.$
As optimization progresses, the increasing sharpness $s_{t}$ ensures that constraint satisfaction is gradually enforced, allowing initial placement flexibility while converging smoothly to the final aligned or evenly spaced configuration. Note that this projection is not performed after the optimizer descent step. Otherwise, it is not aware of this projection. The projection is applied after the tentative descent for step size estimation in the Nesterov optimizer (see Alg. 1 Line 18).
@@ 3.5. Stabilizing Mixed-Size PIC Placement via Blockwise Adaptive Nesterov Optimizer
As emphasized in Section 3.1, the significant heterogeneity in component sizes presents a major challenge for stability and convergence in PIC placement. To address this, we introduce a Blockwise-Adaptive Barzilai–Borwein (BBB) step size scheme in Nesterov-accelerated gradient descent optimizer in Alg. 1, named BNAG, specifically tailored for mixed-size PIC layouts. In our formulation, all placement variables are partitioned into 4 logical blocks: movable instances and dummy fillers in both the $x$ and $y$ dimensions. An independent Barzilai–Borwein (3) step size is computed for each block, and the resulting values are truncated to prevent instability. This blockwise treatment allows us to stabilize the motion of large nodes by decoupling their updates from smaller cells and encourages faster convergence of filler cells to surround movable instances and fill their spacing. While the Barzilai–Borwein (or inverse Lipschitz) step size is effective for acceleration, it often overestimates the appropriate step size, especially when certain cells are already near-optimal. To mitigate this and ensure convergence, we apply a global cosine annealing schedule to gradually reduce the effective step size over time. This helps maintain robustness in the presence of highly uneven cell dimensions and density. Additionally, we project $v_{k+1}$ onto the feasible placement region to correct for potential extrapolation beyond legal bounds in Nesterov’s momentum. This guarantees that gradient evaluations occur at valid locations, preserving placement legality and improving stability.
@@
[…]
@@
[…]